### v2

# Mavors: Multi-granularity Video Representation for Multimodal Large Language Model

Abstract
Long-context video understanding in multimodal large language models (MLLMs) faces a critical challenge: balancing computational efficiency with the retention of fine-grained spatio-temporal patterns. Existing approaches (e.g., sparse sampling, dense sampling with low resolution, and token compression) suffer from significant information loss in temporal dynamics, spatial details, or subtle interactions, particularly in videos with complex motion or varying resolutions. To address this, we propose Mavors, a novel framework that introduces Multi-granularity video representation for holistic long-video modeling. Specifically, Mavors directly encodes raw video content into latent representations through two core components: 1) an Intra-chunk Vision Encoder (IVE) that preserves high-resolution spatial features via 3D convolutions and Vision Transformers, and 2) an Inter-chunk Feature Aggregator (IFA) that establishes temporal coherence across chunks using transformer-based dependency modeling with chunk-level rotary position encodings. Moreover, the framework unifies image and video understanding by treating images as single-frame videos via sub-image decomposition. Experiments across diverse benchmarks demonstrate Mavors’ superiority in maintaining both spatial fidelity and temporal continuity, significantly outperforming existing methods in tasks requiring fine-grained spatio-temporal reasoning.
1 Introduction
Figure 1: (a) Sparse sampling, which remains the high resolution but loses many details in the unsampled frames; (b) Dense sampling with low resolution, which understands the videos from a large number of frames but would confuse on the low-resolution content; (c) Dense sampling with token compression, which keeps the key tokens on the main characters but suffers from hallucinations owing to the missing of visual tokens; (d) Our Mavors, balancing the demands of resolution and number of frames. Though all these approaches could perform similarly on Video-MME, Mavors significantly improves the caption capability on complex scenes. Note that the words in red and green denote incorrect and correct details, respectively.
Long-context video modeling stands as one of the most crucial capabilities within MLLMs [fuyu-8b, zhang2025videollama3, videochat-flash, Maaz2023VideoChatGPT]. This capability empowers MLLMs to proficiently manage hours-long movies, documentaries, and online video streams, all of which demand sophisticated long video processing. Recent advances in MLLMs perform well in short video understanding. However, it remains challenging to build MLLMs for processing extremely long videos (lasting for hours or even longer). The difficulty lies in how to enable MLLMs to efficiently understand the extremely long video context brought by long videos.
As shown in Figure 1, we have compared three mainstream types of video MLLMs with our method, and provided the video caption results of different methods for better illustration. Specifically, in Figure 1(a), these methods (e.g., LLaVA-Video [zhang2024video], InternVL 2.5 [chen2024expanding]) usually employ the sparse sampling strategy to decrease the number of frames and reduce the computation costs. However, these methods have a significant limitation, where many temporal contexts are lost as many frames are not sampled. Thus, the performance results of video-related tasks, which require detailed temporal contexts from many frames, are degraded a lot for these methods. When compared to methods in Figure 1(a), some methods (e.g., Oryx [liu2024oryx], Qwen2VL [Qwen2VL]) have introduced the strategy of dense sampling with low-resolution input in Figure 1(b). However, for these methods, many spatial contexts are lost as only the low-resolution frames are given, which also significantly degrade the results of video-related tasks requiring detailed spatial contexts, e.g., video captioning. Recently, in Figure 1(c), several works (e.g., VideoLLaMA 3 [zhang2025videollama3], VideoChat-Flash [videochat-flash]) have proposed token compression strategies (e.g., token merge or token dropping), which reduces tokens based on vector or pixel similarity and effectively preserves spatial-temporal features of large visual elements. However, token compression inevitably leads to the loss of information regarding small spatial objects, subtle temporal motions, and interactions among multiple objects, thereby posing challenges for understanding complex scenes.
Therefore, the fundamental problem of video understanding is that existing methods often rely on sparse sampling or token compression strategies and struggle to balance computational efficiency with the retention of fine-grained spatio-temporal patterns, particularly in videos with variable motion, aspect ratios, or resolutions.
To address this problem, as shown in Figure 1(d), we introduce the Mavors method to extract the Multi-granularity video representation for MLLMs. which is designed to process raw video content holistically while preserving both spatial fidelity and temporal coherence. Specifically, Mavors eliminates the information loss inherent in conventional frame sampling or token compression methods by directly encoding consecutive video chunks into latent representations. This approach leverages a two-tier architecture: an Intra-chunk Vision Encoder (IVE) extracts high-resolution spatial features from localized video segments using 3D convolutions and Vision Transformer (ViT) layers, while an Inter-chunk Feature Aggregator (IFA) employs temporal transformer and chunk-level rotary position embeddings (C-RoPE) to model temporal dependencies across chunks. Besides, Mavors further unifies image and video understanding by treating images as single-frame videos by employing a sub-image divide-and-conquer approach for image processing. Moreover, following the common training strategy, we also adopt a multi-stage training paradigm, which includes the modality alignment, temporal understanding enhancement, instruction tuning and DPO training stages.
The contributions of Mavors are shown as follows:
  • We propose the Mavors by utilizing the Multi-granularity video representation for multimodal large language model, which aims to better preserve the spatio-temporal contexts based on dense sampling with chunk modeling.
  • Mavors includes two modules: Intra-chunk Vision Encoder (IVE) and Inter-chunk Feature Aggregator (IFA). IFA encodes consecutive video chunks into latent representation based on 3D convolutions and ViT, and IFA builds the temporal coherence based on the temporal transformer and chunk-level rotary-encoding strategies.
  • Comprehensive experimental results and detailed analysis show the effectiveness and efficiency of Mavors.
2 Related Works
2.1 MLLM Architecture
Current MLLMs employ two architectural strategies for visual processing. The first paradigm is based on cross-attention approach, which maintains frozen model parameters while establishing dynamic visual-language interactions through attention mechanisms [Alayrac2022FlamingoAV] . Alternatively, the second paradigm processes visual content through pretrained encoders (CLIP [radford2021clip], SigLIP [siglip]) before concatenating image tokens with text embeddings for unified language model processing [lin2023vila, liu2023llava, liu2023llava15, li2024llava, liu2024llavanext]. The second paradigm can be readily extensible to video analysis through sequential frame processing [2023videochat, zhang2025videollama3], and many architectural innovations for temporal modeling have been proposed [Jiang2025TokenEfficientLV, wang2025internvideo2, kangaroogroup].
2.2 MLLM for Video Understanding
Existing MLLMs have revealed divergent capabilities in temporal comprehension across different video durations. While existing systems demonstrate proficiency in minute-scale video analysis [2023videochat, videochat-flash, lin2023video], emerging efforts targeting hour-level sequences [gemini, wang2024longllava] face fundamental challenges. To address the challenges of long video modeling, current approaches primarily pursue two optimization directions: (1) context window expansion for large language models [gemini, wang2024longllava, Zhang2024LongCT, longvila] and (2) efficient token compression via spatial-temporal feature distillation [videoccam, li2024llamavid, shu2024video, song2023moviechat, TanKoala2024, Weng2024LongVLMEL]. For the first strategy, though theoretically enabling long-sequence processing, suffers from impractical computational overhead, which bring significant challenges for practical applications. In contrast, recent token compression methods like LLaMA-VID [li2024llamavid] achieve compression rates at the cost of discarding subtle details, which results in performance degradation on standard video understanding benchmarks. When compared to the existing works, our Mavors can directly process the raw videos to maintain spatial and temporal details well with acceptable computation costs.
3 Method
3.1 Preliminaries
Necessity of Dense Sampling with High Resolution. As shown in Figure 2 and Figure 3, we have compared the results of two popular video MLLMs (i.e., Qwen2.5-VL-7B [qwen2.5vl] and Oryx-1.5-7B [liu2024oryx]) on two representative benchmarks (i.e., Video-MME [videomme] and DREAM-1K [wang2024tarsier]). Specifically, the Video-MME focuses on multiple-choice question answering based on video content and requires a better understanding of the temporal relations between different frames. DREAM-1K involves open-ended video captioning, where models must generate detailed descriptions of the main events in the video. Thus, both the spatial and temporal fine-grained details are important. In Figure 2, we observe that performance increases a lot when increasing the number of frames, which shows the necessity of dense sampling with more frames. In Figure 3, performance results on Video-MME are relatively stable for both MLLMs. For this phenomenon, we assume that understanding fine spatial details is not vital for Video-MME. In contrast, the results on DREAM-1K increase a lot, which demonstrates the necessity of high resolution.
In summary, as real-world video understanding tasks usually rely on understanding the fine-grained spatiotemporal contexts well, it is important to design video MLLMs by sampling dense and high-resolution frames and maintaining efficiency.
(a) Video-MME
(b) Dream1K
Figure 2: The impact of the number of frames (720P).
(a) Video-MME
(b) Dream1K
Figure 3: The impact of the resolution of frames (64 frames).
Figure 4: The architecture of Mavors.
3.2 Overview of Mavors
In Figure 4, the key objective of Mavors is to enhance the video understanding capability by introducing an efficient video encoding strategy based on dense sampling with high resolution strategy.
Specifically, Mavors employs a video encoder that directly processes pixel information from video chunks, converting them into latent representations. Figure 4 illustrates the overview of Mavors when dealing with video content and images. We consider an input video $S_{\text{V}}\in\mathbb{R}^{W_{\text{V}}\times H_{\text{V}}\times 3\times T_{\text{V}}}$ or an image $S_{\text{I}}\in\mathbb{R}^{W_{\text{I}}\times H_{\text{I}}\times 3}$, where $W_{\text{V}},H_{\text{V}}$ and $W_{\text{I}},H_{\text{I}}$ denote the respective widths and heights, and $T_{\text{V}}$ denotes the total number of video frames. Mavors follows the auto-regressive architecture to generate a textual response based on a given textual instruction. Specifically, in Mavors, we first perform the preprocessing on the raw videos or images to obtain the model input. Then, we employ an intra-chunk vision encoder and an inter-chunk feature aggregator to fully comprehend videos, so that the spatial and temporal details would be remained. Following the mainstream architecture of MLLMs, the temporally integrated features are passed through an MLP projector for modality alignment before being input to the LLM.
3.3 Intra-chunk Vision Encoder
Mavors partitions the video frames into $c_{\text{V}}=\lceil\frac{T_{\text{V}}}{F}\rceil$ video chunks, where each chunk contains $F$ consecutive frames describing the dynamic scenes and temporal events, i.e., $C_{1,\ldots,c_{\text{V}}}=\text{Partition}(S_{\text{V}})$. Intra-chunk vision encoder is designed to represent the vision features of the video content. It begins with 3D convolutions applied to individual video chunks, and we would obtain the visual feature $\mathcal{F}_{i}$ for the $i$-th chunk as follows:
$\mathcal{F}_{i}=\text{Conv}(C_{i})/F\in\mathbb{R}^{n_{\text{V}}\times d_{\text{V}}},i=1,\ldots,c_{\text{V}},$ (1)
where $n_{\text{V}}$ indicates the number of visual features per video chunk, and $d_{\text{V}}$ denotes the dimension of the visual features. We then adopt a standard ViT with parameter $\theta_{\text{ViT}}$ to capture high-level spatial-temporal features, denoted as $\hat{\mathcal{H}}_{i}$, within the $i$-th chunk. To manage the computational load and complexity for the downstream LLM module arising from a large number of tokens, we apply a 2x2 pooling layer on $\hat{\mathcal{H}}_{i}$ to obtain $\mathcal{H}_{i}\in\mathbb{R}^{n_{\text{V}}/4\times d_{\text{V}}}$.
We initialize $\theta_{\text{ViT}}$ by SigLIP weights. Specifically, the 2D convolutional kernels from SigLIP are replicated $F$ times along the temporal dimension to form the 3D kernels. As the resulting visual features are divided by $F$ in Eqn. (1), the spatial absolute position embedding is added to the feature vectors towards the corresponding pixel patches. This ensures that the model’s initial behavior precisely matches its capability for single image-text understanding.
3.4 Inter-chunk Feature Aggregator
The intra-chunk vision encoder mainly captures the high-level visual features within video chunks. Mavors leverages the the inter-chunk feature aggregator, to integrate temporal information across the multiple video chunks of the complete video. First, we concatenate the high-level visual features to form the original feature sequence as follows:
$\chi^{(0)}=\text{Concat}(\mathcal{H}_{1,\ldots,c_{\text{V}}}).$ (2)
Inter-chunk feature aggregator consists of $L_{\text{inter}}$ Transformer layers with Causal Attention. To identify the sequential order of the visual features, we propose chunk-level Rotary Encoding (C-RoPE) to the Transformer layers, so that the temporal information can be correctly retained. Specifically, the causal scaled dot product (SDP) attention in the $j$-th Transformer layer would be calculated by
$\mathcal{Q}_{\text{Inter}}^{(j)},\mathcal{K}_{\text{Inter}}^{(j)},\mathcal{V}_{\text{Inter}}^{(j)}=\text{Linear}(\chi^{(j-1)}),$ (3)
$\text{SDP}(q_{\iota}^{(j)},k_{\iota^{\prime}}^{(j)})$ $=\text{C-RoPE}(q_{\iota}^{(j)},k_{\iota^{\prime}}^{(j)};\lceil\frac{4\iota}{n_{\text{V}}}\rceil,\lceil\frac{4\iota^{\prime}}{n_{\text{V}}}\rceil)$ (4)
$=q_{\iota}^{(j)}R_{\lfloor\frac{4\iota}{n_{\text{V}}}\rfloor-\lfloor\frac{4\iota^{\prime}}{n_{\text{V}}}\rfloor}k_{\iota^{\prime}}^{(j)\intercal},$
$\forall q_{\iota}^{(j)}\in\mathcal{Q}_{\text{Inter}}^{(j)},k_{\iota^{\prime}}^{(j)}\in\mathcal{K}_{\text{Inter}}^{(j)}$
Here, $R$ represents the rotation matrix. In practice, we would transcode the video into fixed FPS, so that the index of the video chunk can be identified from the actual timestamp of the first frame of the chunk. In the remaining process of the Transformer layer, we follow
$\mu^{{j}}=\text{softmax}(\text{SDP}(\mathcal{Q}_{\text{Inter}}^{(j)},\mathcal{K}_{\text{Inter}}^{(j)})),$ (5)
$\chi^{(j)}=\mu^{{j}}\mathcal{V}_{\text{Inter}}^{(j)}.$ (6)
We then feed $\chi^{(L_{\text{Inter}})}$ to the MLP projector to obtain the visual tokens, where the feature dimension of these visual tokens is the same as the feature dimension of textual tokens in LLM.
Figure 5: The dynamic resolution strategy in Mavors.
Figure 6: Training paradigm of different stages.
3.5 Preprocessing
Video Preprocessing. The video processing strategy of Mavors varies based on the video length. Specifically, videos with short lengths are directly processed into chunks. To accommodate long videos, we employ an initial step of accelerated playback achieved through frame dropping, thereby reducing the total frame count to be compatible with Mavors processing limits. Specifically, the position IDs utilized by C-RoPE correspond to timestamps derived from the original, non-accelerated video timeline. This mechanism informs the model that the processed frames are not temporally contiguous. While alternative strategies for very long video comprehension exist, e.g., in-video Retrieval-Augmented Generation (RAG) [luo2024video], they represent an orthogonal direction to Mavors.
Meanwhile, Mavors could process videos with arbitrary resolutions and aspect ratios. Specifically, Mavors employs a dynamic resolution strategy to maintain the original aspect ratio of the video frames, avoiding distortion artifacts that can arise from fixed-shape resizing. The resized video frames roughly keep the original aspect ratio and match the number of pixels in the ViT’s pretraining images. For example, given the frames with the $(W_{\text{V}},H_{\text{V}})$ resolution and the ViT’s pretrained image resolution $(R_{v},R_{v})$, Mavors will rescale the frames into the resolution of $(R_{v}*\sqrt{W_{\text{V}}/H_{\text{V}}},R_{v}*\sqrt{H_{\text{V}}/W_{\text{V}}})$. We also resize the positional embedding of patches, following SigLIP [siglip]. Specifically, the positional embedding of the video chunk in the $(x,y)$ position, denoted as $E(x,y)$, will be formulated as:
$E(x,y)=E_{v}(x*P_{v}/P_{W},y*(P_{v}/P_{H})),$ (7)
where $(P_{W},P_{H})$ is the number of patches in the video chunk. $P_{v}$ and $E_{v}(x,y)$ are the number of patches and the positional embedding in the ViT’s pretraining images, respectively.
Image Preprocessing. As shown in Figure 5, Mavors first partitions the raw image into several sub-images, and then leverages the thumbnail of the original image and all sub-images into the vision encoder. Besides, Mavors incorporates a special design in the feature aggregator to accommodate the joint training of videos and images. The details are as follows.
First, as image understanding tasks often require spatial details, we follow the image partition method in [yao2024minicpmvgpt4vlevelmllm] and support dynamic resolution for processing high-resolution images, where the raw image will be partitioned into multiple sub-images and the size of these sub-images is supposed to match the number of pixels in the ViT’s pretraining. Specifically, we first determine the ideal number of sub-images $N_{s}=\left\lfloor(W_{\text{I}}\times H_{\text{I}})/R_{v}^{2}\right\rfloor$, where ($W_{\text{I}},H_{\text{I}}$) is the resolution of the original raw image and ($R_{v},R_{v}$) is the resolution of the ViT’s pretraining images. Next, we identify potential partition configurations by finding pairs of integers $(m,n)$, representing the number of columns and rows, respectively, such that their product equals the target number of slices $N_{s}$. These pairs form the set $\mathcal{C}_{N_{s}}=\{(m,n)|m\times n=N_{s},m,n\in\mathbb{Z}\}$. Then, we select the best configuration $(m^{*},n^{*})$ from $\tilde{C}=\mathcal{C}_{N_{s}-1}\cup\mathcal{C}_{N_{s}}\cup\mathcal{C}_{N_{s}+1}$ based on the following criteria:
$\quad(m^{*},n^{*})=\arg\min_{(m,n)\in\tilde{C}}\left|\log\frac{W_{\text{I}}}{H_{\text{I}}}-\log\frac{m}{n}\right|.$ (8)
We will leverage the thumbnail of the original raw image $I_{0}$ and all sub-images $I_{1},...,I_{m^{*}\times n^{*}}$ as the input of the vision encoder. Before feeding into the vision encoder, we will rescale the original image and the sub-images, which have more pixels than the ViT’s pretraining images. We use the same dynamic resolution strategy as video processing.
Second, when compared to video processing, the feature aggregator operates on the features extracted from each sub-image independently, thus avoiding redundant temporal relationships. Furthermore, given that the model must process both images and videos, the representation of an image (treated as a single frame) is replicated across all temporal positions within the input sequence. Placing the image representation at only a single temporal position would cause the model parameters to become biased towards that static position, ultimately hindering the model’s capacity to perceive temporal information effectively in video sequences.
4 Training Paradigm
In Figure 6, multi-stage training is adopted, serving to improve the collaboration of the video encoder and LLM and the performance of multimodal tasks. Given SigLIP’s robust image understanding performance, we forgo an independent CLIP training phase to avoid redundancy. Instead, we adopt a tailored initialization strategy to ensure compatibility with both video and image inputs, where the 2D convolutional kernels from SigLIP are replicated $F$ times along the temporal dimension to form the 3D kernels. Then, we leverage multiple training stages to progressively build a vision encoder that maintains image understanding while effectively encoding spatio-temporal information of videos. The data used for training Mavors is detailed in Appendix A.
Stage 1: Modality Alignment. As SigLIP’s training involved alignment with the T5 model [2020t5], the first stage aims to align the semantic space of the vision encoder with the LLM’s semantic space. In this stage, we train the inter-chunk feature aggregator and the MLP projector, while keeping the LLM and the intra-chunk vision encoder frozen. Although the model exhibits only coarse video comprehension at this stage, the principal aim is to achieve modality alignment and instill basic temporal understanding. Therefore, we prioritize diverse, general-concept image-text pairs and short video-text pairs with low complexity (e.g., LAION [laion400m] and PANDA-70M[chen2024panda]), thereby avoiding excessively difficult data that could impede the development of foundational abilities.
Stage 1.5: Temporal Understanding Enhancement. Subsequent to Stage 1, we implement Stage 1.5, which focuses on enhancing the video encoder’s capacity for genuine video comprehension. Based on the modality alignment from Stage 1, parameter updates are performed on all components excluding the LLM. For data selection in this stage, we augment the initial dataset with standard computer vision (CV) tasks applied to images and short video chunks, such as captioning, classification, OCR, interleaved image-text, and perception QA.
Stage 2: Multitask Instruction Tuning. In Stage 2, the primary objective is to adapt the model for a range of multi-modal tasks, leveraging data formats including text-only, single-image, multi-images, and complex video. Beyond standard CV tasks, we incorporate grounding tasks and temporal grounding tasks to enhance the model’s perception of spatio-temporal details. Similar to the practice in Qwen2.5VL [qwen2.5vl], we find that representing bounding boxes using plain text coordinates yields performance comparable to using special tokens; consequently, we adopt the plain text representation. This stage also activates the sub-image partitioning paradigm to enhance the model’s image understanding capabilities. All model parameters are unfrozen and trained on a large dataset, allowing for extensive self-adjustment. Upon completion, the model possesses significant world knowledge, semantic understanding, and logical reasoning abilities, though its application is initially limited by the specific tasks and query formats encountered. Therefore, towards the end of this stage, we introduce more diverse data types, covering a broader spectrum of real-world task scenarios and textual query formulations.
Stage 3: DPO Training. Our empirical evaluations reveal that while the previously described training procedure yields strong leaderboard performance, the resulting model exhibits distinct patterns. Specifically, for QA tasks, the model tends to generate overly concise responses, likely due to extensive training on multiple-choice or short-answer datasets. Conversely, for descriptive tasks, the model fails to terminate generation appropriately. To mitigate these issues, we incorporate a Direct Preference Optimization (DPO) [Rafailov2023DirectPO] stage following Stage 2. The preference dataset mainly covers three domains: open-ended QA, image captioning, and video captioning. More details can be found in Appendix A.
Loss Function. We employ the next-token-prediction (NTP) training methodology in all training stages except the DPO stage. During DPO training, we employ the standard DPO loss.
5 Experiments
Model Size MMWorld PerceptionTest Video-MME MLVU MVBench EventHallusion TempCompass VinoGround DREAM-1K
GPT-4o-20240806 - 62.5 - 71.9 64.6 64.6 92.0 73.8 38.9 39.2
Gemini-1.5-Pro - - - 75.0 - 60.5 80.3 67.1 22.9 36.2
LLaVA-OneVision 7B 59.2 56.9 58.9 64.8 56.7 64.3 61.4 26.2 31.9
InternVL 2.5 8B 62.2 65.0 64.3 67.0 72.0 64.1 71.4 24.0 29.7
NVILA 8B 55.2 55.5 64.2 70.1 68.1 69.9 66.5 20.2 26.9
LLaVA-Video 7B 60.1 67.5 63.6 67.2 58.6 70.7 65.7 26.9 33.3
Oryx-1.5 7B 58.8 70.3 59.0 63.8 67.5 61.3 60.2 22.3 32.5
Qwen2.5-VL 7B 61.3 66.2 65.1 70.2 69.6 66.5 71.4 34.6 32.6
VideoLLaMA3 7B 56.4 72.8 66.2 73.0 69.7 63.4 68.1 31.3 30.5
VideoChat-Flash 7B 57.9 74.7 65.3 74.7 74.0 66.4 70.0 33.3 29.5
Slow-fast MLLM 7B 58.2 69.7 60.2 60.4 68.9 67.4 69.9 27.1 33.2
Qwen2.5-VL 72B 73.1 73.2 73.3 76.6 70.4 76.3 79.1 58.6 35.1
InternVL 2.5 78B 77.2 73.5 72.1 76.6 76.4 67.7 75.5 38.7 30.3
Mavors (Ours) 7B 68.1 70.3 65.0 69.8 68.0 73.5 77.4 36.9 39.4
Table 1: Performance on video benchmarks. Most of the scores are from their original studies. The others are reproduced following the official benchmark recommendation.
5.1 Experimental Setup
Implementation Details. The Mavors model utilizes Qwen2.5-7B as its language model module, with the intra-chunk vision encoder initialized using SigLIP weights. To balance effectiveness and efficiency, the frame count per video chunk, $F$, is set to 16. The inter-chunk feature aggregator consists of $L_{\text{Inter}}$=3 layers. The training is conducted on 416 GPUs. Given the model’s moderate size, we employed DeepSpeed with ZeRO stage 2 optimization. As mentioned in Section 4, the pre-training proceeded in three stages: Stage 1 used approximately 127 million samples with a global batch size of 6,656, taking 71 hours; Stage 1.5 used 52 million samples with a global batch size of 3,328, taking 177 hours; and Stage 2 used 19 million samples with a global batch size of 1,664, requiring 28 hours. The learning rates for the LLM and projector are set to 1e-5 in both Stage 1 and Stage 1.5, with a constant learning rate schedule applied during these phases. In Stage 2 and DPO, the learning rate was initialized at the same value (1e-5) as the preceding stages but followed a cosine decay schedule, gradually reducing to 1/10th of its initial value. Meanwhile, the learning rates for the inter-chunk feature aggregator and intra-chunk vision encoder remained fixed at 1/10th of the LLM’s learning rate across all training stages.
For inference, Mavors is adapted using the vLLM framework [kwon2023efficient]. Since Mavors requires comprehensive video encoding and frame preprocessing occurs on the CPU, the CPU processor can thus become a bottleneck. Recognizing that the intra-chunk vision encoder’s computation is a one-time GPU operation per video, with results stored in the LLM’s KV cache, we overlaps the pipeline. Specifically, the intra-chunk vision encoder and inter-chunk feature aggregator execute directly on the GPU, while the language model component leverages vLLM. This separation can effectively balance CPU-bound preprocessing, compute-intensive visual encoding (Intra/Inter), and language model inference. More details of the inference efficiency can be found in Appendix B.
Baseline Models. We select several representative video models for performance comparison. We include GPT-4o-20240806 [hurst2024gpt] and Gemini-1.5-Pro-002 [gemini] as the closed-source APIs baselines. Standard auto-regressive models using resolution-preserving frame sampling are represented by LLaVA-OneVision [li2024llava] and InternVL 2.5 [chen2024expanding]. For video understanding tasks, we add models based on: (a) high-performing sparse frame sampling (NVILA [liu2024nvila], LLaVA-Video [zhang2024video]); (b) dense sampling with lower resolution (Qwen2.5-VL [qwen2.5vl], Oryx-1.5 [liu2024oryx]); (c) dense sampling with token compression (VideoChat-Flash [videochat-flash], VideoLLaMA3 [zhang2025videollama3]); and (d) slow-fast architecture, a special frame sampling strategy (Slow-fast MLLM [shi2025slow]). Regarding image tasks, as some video-centric models either lack image input (e.g., VideoChat-Flash) or are not SOTA on image tasks, we include four strong models on QA/Caption benchmarks: GLM-4V [wang2024cogvlm], Qwen2.5-VL, DeepSeek-VL2 [deepseek_vl2] and CogVLM2 [hong2024cogvlm2]. Crucially, aside from prompt modifications, no benchmark-specific hyperparameters (e.g., frame sampling, resolution) were tuned during evaluation for any model, including Mavors.
Benchmarks. Video understanding capabilities are assessed across general knowledge QA (MMWorld [hemmworld], PerceptionTest [patraucean2023perception]), long-video QA (Video-MME [videomme], MLVU [MLVU]), event understanding QA (MVBench [li2024mvbench], EventHallusion [zhang2024eventhallusion]), temporal understanding QA (TempCompass [tempcompass], VinoGround [vinoground]), and captioning (DREAM-1K [wang2024tarsier]). Image understanding evaluation includes comprehensive capabilities (MMMU [yue2023mmmu]), cognitive understanding (MathVista [mathvista], AI2D [kembhavi2016diagram]), and captioning (CapsBench [liu2024playground]). More experiment details can be found in Appendix C.
5.2 Main Results
Video Understanding. Table 1 presents a performance comparison of Mavors against baseline models on various video benchmarks. Approaches employing dense frame sampling with lower resolution demonstrate strong performance on long video QA by incorporating extensive temporal information, but exhibit limitations in understanding spatial details for knowledge-intensive and captioning tasks. token compression strategies show a similar pattern, yielding excellent scores on long video QA due to abundant temporal cues, but their merging of non-primary tokens compromises the comprehension of environmental context, resulting in marked deficiencies, especially in captioning. In contrast, sparse frame sampling approaches, which inherently lose temporal detail and consequently perform less effectively on event understanding QA. Mavors’s multi-granularity video understanding framework successfully balances these trade-offs. Leveraging efficient visual information compression, Mavors delivers performance on long video QA nearly on par with dense sampling and token compression techniques, while preserving robust capabilities for knowledge-based and temporal reasoning tasks, eliminating the need for dataset-specific hyperparameter tuning. The substantial gains observed for Mavors in captioning highlight the effectiveness in achieving accurate and comprehensive understanding of entire video events.
Image Understanding. Table 2 compares Mavors’s performance against baseline models on image benchmarks. Mavors achieves performance on par with similarly-sized image understanding models in Image QA. Its captioning performance is particularly strong, surpassing even 72B models. This effectiveness is partly due to Mavors’s architecture: images and videos offer complementary visual perception within the intra-chunk vision encoder, yet are processed without mutual interference by the inter-chunk feature aggregator.
Model Size MMMU MathVista AI2D CapsBench
GPT-4o-20240806 - 69.9 62.9 84.7 67.3
Gemini-1.5-Pro - 60.6 58.3 79.1 71.2
CogVLM2 8B 42.6 38.7 73.4 50.9
GLM-4V 9B 46.9 52.2 71.2 61.0
LLaVA-OneVision 7B 47.9 62.6 82.4 57.4
InternVL 2.5 8B 56.2 64.5 84.6 66.5
Qwen2.5-VL 7B 58.0 68.1 84.3 64.9
DeepSeek-VL2 27B 54.0 63.9 83.8 61.3
Qwen2.5-VL 72B 68.2 74.2 88.5 70.1
InternVL 2.5 78B 70.0 70.6 89.1 68.5
Mavors (Ours) 7B 53.2 69.2 84.3 75.2
Table 2: Performance on image benchmarks.
Figure 7: Performance with different numbers of frames in a video chunk.
Figure 8: Performance with different token compression ratios.
Figure 9: The dynamic of training losses across different stages for Mavors.
5.3 Ablation Studies
We conduct a series of ablation studies to validate our model design. Given the extensive training time required for the full training paradigm, these ablations utilize standard compositive datasets and train various versions up to the completion of Stage 2. Specifically, Stage 1 employs LLaVA-Pretrain-558K [liu2023llava15] and LLaVA-Hound-Pretrain [llava-hound]; Stage 1.5 uses M4-Instruct [li2024llavanextinterleave] and ShareGPT4o [cui2025comprehensive]; and Stage 2 utilizes LLaVA-OneVision and LLaVA-Video. This approach reduces the duration of a full training cycle to under 24 hours with 64 GPUs. Performance is subsequently monitored using MMMU, MathVista, and CapsBench for image understanding capabilities, and Video-MME, Vinoground, and DREAM-1K for video understanding capabilities.
Effect of the Number of Frames in a Video Chunk. We conduct experiments with four settings, varying a parameter $F$ with values of 4, 8, 16, and 32. Upon the preliminary study evaluating video captioning performance on the validation set of KVQ [lu2024kvq], we observe that configurations with $F=8$ or $F=16$ yield more accurate and comprehensive captions. To ensure exposure to richer visual information, we finalize the $F=16$ setting. We further evaluate these four model variants on six benchmark datasets in Figure 9. On image-based tasks, we observe a marginal improvement in performance metrics with increasing $F$. We hypothesize that this improvement stems from the model’s increased exposure to individual frames during video processing when $F$ is larger, thereby enhancing its image understanding capabilities. Conversely, for video understanding tasks, performance degrades significantly for $F=4$ due to insufficient temporal information and for $F=32$, likely due to excessive information compression.
$L_{\text{Inter}}$ MMMU MathVista CapsBench Video-MME VinoGround DREAM-1K
0 50.3 63.0 51.4 61.0 27.9 30.2
1 51.5 63.3 50.6 60.9 30.6 32.4
3 52.0 62.6 50.6 61.1 31.1 33.8
5 49.8 61.9 50.3 61.1 31.2 33.6
Table 3: Ablation on layers of Transformers in IFA.
Effect of the IFA Module. We establish two baseline models for comparison in Table 3. The first baseline completely removes the inter-chunk feature aggregator ($L_{\text{Inter}}$=0), where the output from the IVE module is passed directly through a projector and then concatenated with the LLM’s input sequence. In this setup, the integration of temporal and spatial information relies solely on the LLM. The second baseline utilizes only a single Transformer layer ($L_{\text{Inter}}$=1) for the aggregator, thereby reducing its computational complexity. In Table 3, on image evaluation tasks, removing the Transformer ($L_{\text{Inter}}$=0) shows a slight advantage, potentially due to the lower parameter count facilitating faster convergence on static perception tasks. However, for video evaluation, we observe that a deeper inter-chunk feature aggregator ($L_{\text{Inter}}$=3) enhances the model’s understanding, leading to better scores, although with diminishing marginal returns. Considering model complexity and convergence difficulty, $L_{\text{Inter}}$=3 should be an efficient configuration of Mavors.
RoPE MMMU MathVista CapsBench Video-MME VinoGround DREAM-1K
Standard 51.9 62.6 50.7 61.0 30.3 32.9
C-RoPE 52.0 62.6 50.6 61.1 31.1 33.8
(+0.1) (+0.0) (-0.1) (+0.1) (+0.8) (+0.9)
Table 4: Ablation on C-RoPE.
Effect of C-RoPE. To assess the performance of C-RoPE, we replace it with the standard RoPE implementation and monitor changes in the Mavors model’s visual understanding performance. Table 4 shows the performance across six metrics. For image understanding, given that the IFA architecture processes sub-images independently, both RoPE variants perform comparably. Conversely, for video understanding, C-RoPE outperforms standard RoPE by an average of 0.6 points. It indicates that standard RoPE suffers from differentiating intra-chunk from inter-chunk tokens and may hinder temporal sequence modeling. These findings demonstrate the efficacy and importance of C-RoPE within the IFA architecture.
5.4 Further Analysis
Analysis on the Ratios of Token Compression. We apply token compression techniques within Mavors to decrease the number of tokens on each video chunk. Specifically, prior to the inter-chunk feature aggregator, we compute similarity between features at corresponding indices in adjacent chunks. Tokens exceeding a predefined similarity threshold are merged via averaging, retaining the positional ID from the earlier chunk. We vary thresholds to achieve different token reduction ratios, summarized in Figure 9. Results indicate that Mavors’ performance on video QA remains largely unaffected with token reductions up to 60%. Conversely, a significant performance degradation is observed for video captioning. This suggests that token compression on Mavors can be a feasible strategy for reducing inference costs in long-video QA applications. We provide two representative cases in Appendix F.
Stage MMMU CapsBench Video-MME DREAM-1K
Stage 1 36.3 54.8 48.4 23.6
Stage 1.5 47.3 62.5 53.9 26.3
Stage 2 53.0 73.4 65.0 38.9
DPO 53.2 75.2 65.0 39.2
Table 5: Results of different training stages.
Analysis on the Training Dynamics. Table 5 shows the results on the image QA dataset (MMMU), image caption dataset (CapsBench), video QA dataset (Video-MME) and video caption dataset (DREAM-1K) at different stages. The performance on all four datasets improves consistently across the three training stages (Stage 1, Stage 1.5, and Stage 2), indicating that each stage contributes positively to the model’s ability to handle different tasks and modalities. The DPO stage provides further improvements. Note that we also provide the training loss curve of Mavors in Figure 9.
Visualization. We pick a complex video cut from DREAM-1K and present the captions generated by Qwen2.5VL-7B and Mavors-7B in Figure 10. Despite processing densely sampled frames, Qwen2.5VL-7B fails to capture many details (e.g., omitting the mention of a cow driving), leading to flawed inferences (words in red). In contrast, we observe that Mavors-7B predicts fine-grained and correct details (words in greed), which show the effect of our Mavors-7B.
Figure 10: Comparison of generated video captions from Qwen2.5-VL-7B and Mavors-7B.
6 Conclusion
In this work, we present Mavors, a novel framework for holistic long-context video understanding in MLLMs. Mavors introduces multi-granularity video representation based on Intra-chunk Vision Encoder (IVE) and Inter-chunk Feature Aggregator (IFA) to preserve both spatial details and temporal dynamics and maintain high efficiency. Extensive experiments on multiple benchmarks demonstrate the effectiveness and efficiency of our Mavors.
Appendix
Task Dataset
Stage 1 Datasets
Image Caption LAION (EN 6.7M, ZH 3.2M) [laion400m], Conceptual Captions (7.3M) [sharma2018conceptual], SBU (0.8M) [ordonez2011im2text], COYO (11M) [coyo-700m], WuKong (2.9M) [gu2022wukong], LAION COCO (16M) [lacion-coco], OMEGA Image Caption (79M) [omegaproject]
Video Caption InternVid-10M-FLT (1.6M) [wang2023internvid], Panda-70M (0.9M) [chen2024panda], OMEGA Video Caption (4M) [omegaproject]
Stage 1.5 Datasets
Image Caption Met-meme [xu2022metmeme], PD12M [meyer2024public], dalle3 [dalle3], GBC10M [hsieh2024graph], DenseFusion-1M [li2024densefusion], GameBunny [taesiri2024videogamebunny], MERMAID [toh2023mermaid], CC12M (1M) [changpinyo2021cc12m], BLIP3 [awadalla2024blip3], AllSeeingV2 [wang2024all]
Video Caption ChronoMagic [yuan2024chronomagic], VideoChatGPT [Maaz2023VideoChatGPT], YouCook2 [zhou2018towards], CelebV [yu2023celebv], SthSthV2 [goyal2017something], MiraData [ju2024miradata], Hacs [zhao2019hacs], OpenVid-1M [nan2024openvid], Kinetics_700 [carreira2019short], ShareGPT4Video [chen2024sharegpt4video], Vript [yang2024vript], Shot2Story [han2023shot2story20k], ShareGemini [sharegemini]
Question Answering MMDU [liu2024mmdu], MMiT [monfort2021multi]
Knowledge Wikipedia [wikidump], Wikimedia [wikidump], WIT [srinivasan2021wit]
Code WebSight [laurençon2024unlocking]
OCR LSVT [sun2019icdar], ArT [chng2019icdar2019], DocMatix [laurenccon2024building]
Interleaved OBELICS [laurencon2023obelics], PIN [wang2024pin]
Mixed-Task Dataset MMInstruct [liu2024mminstruct], LVD-2M [xiong2024lvd], MMEvol [luo2024mmevol]
Stage 2 Datasets
Instruction Countix [Dwibedi_2020_CVPR], VideoChat [2023videochat], Videogpt+ [maaz2024videogpt+], Openmathinstruct-2 (2M) [toshniwal2024openmathinstruct], RepCountA [hu2022transrac], Vidgen-1m [tan2024vidgen], CompCap [chen2024compcap], Metamath [yu2023metamath], Llava-Onevision [li2024llava], Anytext (0.3M) [tuo2023anytext], Llava-Video [zhang2024video], S-MiT [monfort2021spoken], LSMDC [rohrbach2017movie], Infinity-MM [gu2024infinity], Mantis [Jiang2024MANTISIM], ShareGPT4V [chen2023sharegpt4v], CinePile [rawal2024cinepile], LLaVA-Hound [llava-hound]
Grounding GRIT [Kosmos2], RefCOCO [kazemzadeh2014referitgame]
Temporal Grounding GroundedVideoLLM [wang2024grounded]
Stage 3 (DPO) Datasets
Open-ended QA Llava-Video [zhang2024video] (10K)
Image Caption Llava-Onevision [li2024llava] (10K), DenseFusion-1M [li2024densefusion] (10K)
Video Caption WebVid [bain2021frozen] (8K), Kinetics_700 [carreira2019short] (8K), OOPS [epstein2019oops] (4K)
Table 6: Summary of the training datasets of different stages.
A Training Datasets
The datasets used for training our model at different stages are shown in Table 6. For a number of large-scale datasets, we have randomly selected a specific number of samples. The count of these samples is also indicated in Table 6.
We have also curated two datasets from the OMEGA project [omegaproject], the OMEGA Image Caption (containing 79M samples) and OMEGA Video Caption (containing 4M samples), by sampling videos and images along with their corresponding titles and captions. These two datasets are utilized in the first stage of our model training.
For certain datasets that either lack captions or only possess low-quality ones, for example, CC12M [changpinyo2021cc12m], CelebV [yu2023celebv], Hacs [zhao2019hacs], and Kinetics_700 [carreira2019short], we carefully designed a pipeline to generate high-quality captions. Initially, we utilized Qwen2VL-72B [Qwen2VL], InternVL2.5-78B-MPO [chen2024expanding] and Tarsier-34B [wang2024tarsier] (video only) to describe these samples in detail. Subsequently, we used DeepSeek-R1-Distill-Llama-70B [deepseekai2025deepseekr1] to amalgamate captions generated by different models while attempting to resolve all inconsistencies using its COT capabilities. The captions produced by this process generally demonstrated superior quality and comprehensibility.
We observed that many composite datasets incorporate content from established standalone datasets, leading to potential data redundancy. To address this, we implemented a deduplication process for identical samples (images or videos). Specifically, we calculated the Perplexity (PPL) of the associated text using the Qwen2VL-72B [Qwen2VL] model, distinguishing between QA and Captioning tasks. For duplicate visual content within QA tasks, we retained the two samples exhibiting the lowest text PPL scores. For Captioning tasks, one sample was randomly selected from the two with the lowest PPL for inclusion in our training set.
For the data in the DPO stage, we selected a specific number of samples from the corresponding datasets. The preference datasets were then generated in accordance with the following methods:
  1. 1.
    Open-ended QA: Positive examples are generated by prompting the model with diverse inputs to produce responses that are correct, of appropriate length, and properly terminated. Negative examples are derived from the same inputs by adjusting the sampling temperature to elicit incorrect or overly brief answers.
  2. 2.
    Image Captioning: Multiple candidate captions are generated per image using the model under high temperatures. These candidates are then ranked according to a predefined scoring strategy, forming positive (higher-ranked) and negative (lower-ranked) pairs for DPO training.
  3. 3.
    Video Captioning: Captions generated from the original video serve as positive examples. Negative examples are created by captioning the video after segmenting it into four equal parts and shuffling their temporal order.
B Analysis on the Inference Costs
Qwen2.5VL-7B Mavors-7B
Images Prefilling (ms) 397 392
Decoding (token/s) 23 30
Videos Prefilling (ms) 1,225 448
Decoding (token/s) 22 30
Table 7: Inference efficiency between Qwen2.5VL-7B and Mavors-7B. Model is better when Prefilling (ms) is lower and Decoding (token/s) is larger.
We evaluate the inference performance of Qwen2.5VL-7B and Mavors-7B using an GPU. Initially, we measure the execution time of the model.generate function via the standard HuggingFace implementation (with FlashAttention-2 enabled) to capture the core model execution time, excluding video preprocessing. Table 7 summarizes the inference times for both models on the DREAM-1K and CapsBench video captioning tasks. The results show that Mavors’ more efficient video representation reduces both the ViT computations and the language model’s context window requirements. Consequently, Mavors-7B demonstrates significant speed improvements on video understanding tasks, achieving 2.7x faster prefill and 1.4x faster decoding compared to Qwen2.5VL-7B. Furthermore, integrating the vLLM inference framework with overlapping vision preprocessing enables 2.5s per image in CapsBench and 3.7s per video in DREAK-1K, reducing from about 13s per image and 20s per video respectively. These findings indicate that Mavors provides an economical solution for scenarios requiring frequent or high-volume multimodal model inference.
C Details of Experiments
Evaluation Setup. To ensure a standardized and reproducible evaluation, we conduct experiments on both open-source and closed-source models using consistent protocols. For open-source models, we adopt the lmms-eval framework [zhang2024lmms], which offers a unified pipeline tailored for benchmarking MLLMs. All open-source models are evaluated using the officially released checkpoints to preserve the integrity of reported results. To maintain experimental stability, we fix the decoding strategy to greedy decoding, set the maximum number of generated tokens to 1024. Image and video resolution, along with other preprocessing settings, follow the default configurations provided by the lmms-evak framework or the respective model implementations. For closed-source models, including Gemini-1.5-Pro-002 [gemini] and GPT-4o-20240806 [hurst2024gpt], we access them through their official APIs. Due to the restricted controllability over decoding parameters, we adopt the default generation settings provided by each platform. For benchmarks requiring GPT-based automatic scoring, such as those involving instruction-following or open-ended generation tasks, we follow the evaluation protocol described in the original benchmark papers or apply the default settings specified by the lmms-eval framework to select the judge model. Specifically, for MathVista [mathvista], we use GPT-4-Turbo (1106) as the judge model. For CapsBench [liu2024playground] and MMMU [yue2023mmmu], we adopt GPT-4o (20240806), while for DREAM-1K [wang2024tarsier], we follow the original benchmark and employ GPT-3.5-Turbo (0125) to perform automatic scoring. These choices align with the evaluation protocols used in the respective benchmark papers, ensuring fair and comparable results across models.
Baseline Models. To comprehensively evaluate the performance of our proposed Mavors-7B, we select a diverse set of baseline models tailored to the specific characteristics of both image and video benchmarks.
For image benchmarks, we compare against two leading proprietary models, GPT-4o [hurst2024gpt] and Gemini-1.5-Pro [gemini]. GPT-4o, developed by OpenAI, is capable of processing text, images, and audio in a unified manner and has demonstrated strong performance in visual reasoning tasks. Gemini, developed by Google DeepMind, similarly integrates multimodal capabilities and excels in scenarios requiring complex cross-modal understanding. We also include a range of high-performing open-source MLLMs in our comparison. These include CogVLM2 [hong2024cogvlm2], a model optimized for visual-language understanding in dynamic contexts; GLM-4V [hong2024cogvlm2], which extends the GLM architecture with strong visual recognition capabilities; LLaVA-OneVision [li2024llava], a widely recognized open-source MLLM that integrates a collection of high-quality multimodal datasets, advanced training strategies, and refined model designs to achieve strong performance across image-based benchmarks; InternVL2.5 [chen2024expanding], which is an advanced MLLM series developed by Shanghai Artificial Intelligence Laboratory. Building upon the architecture of InternVL2, it introduces significant enhancements in training strategies and data quality; DeepSeek-VL2 [deepseek_vl2], an MoE-based model balancing scalability and accuracy; and Qwen2.5-VL [qwen2.5vl], a model that significantly enhance general image recognition capabilities, expanding to a vast array of categories, including plants, animals, landmarks, and various products. It also excels in precise object localization, advanced text recognition, and document parsing.
For video benchmarks, we select four representative categories of baseline models, each exemplifying distinct video processing strategies. The first category includes models that employ sparse frame sampling with high performance, such as NVILA [liu2024nvila] and LLaVA-Video [llava-video-sft], which focus on selecting key frames to reduce computational overhead while maintaining contextual understanding. NVILA, developed by NVIDIA, utilizes a “scale-then-compress” paradigm that first increases spatial and temporal resolutions and then compresses visual tokens, enabling efficient processing of high-resolution images and long videos. LLaVA-Video improves video understanding by introducing a high-quality synthetic dataset, LLaVA-Video-178K [llava-video-sft], specifically designed for video instruction-following tasks. Models like Qwen2.5-VL [qwen2.5vl] and Oryx-1.5 [liu2024oryx] adopt dense frame sampling at lower resolutions to achieve a trade-off between information richness and efficiency (we set at most 768 frames in our experiments). Oryx-1.5 is a unified MLLM designed to flexibly and efficiently handle visual inputs with varying spatial scales and temporal lengths, making it well-suited for processing both high-resolution images and extended video sequences. In addition, we include models such as VideoChat-Flash [videochat-flash] and VideoLLaMA3 [zhang2025videollama3], which apply dense sampling combined with token compression to handle long video sequences effectively (up to 1000 frames in our experiments). VideoChat-Flash leverages this strategy to mitigate the computational overhead introduced by dense sampling, enabling effective handling of long-duration inputs without sacrificing performance. Similarly, VideoLLaMA3 integrates token compression with dense sampling to reduce input redundancy, thereby enhancing the model’s ability to understand and reason over extended video content. Finally, we include Slow-fast MLLM [shi2025slow], which employs a specialized dual-pathway sampling mechanism to capture temporal dynamics at multiple granularities. By processing visual inputs through both slow and fast pathways, the model effectively models temporal variations across different timescales.
Benchmarks. It is crucial to comprehensively and objectively assess a model’s capabilities across various aspects and dimensions. To this end, we include a broad range of representative image and video benchmarks in our evaluation.
We adopt MMMU [yue2023mmmu], MathVista [mathvista], AI2D [kembhavi2016diagram], and CapsBench [liu2024playground] as representative image benchmarks, covering a broad range of visual understanding and reasoning tasks.
  • MMMU targets expert-level multimodal reasoning across diverse academic domains, featuring varied visual inputs such as charts, diagrams, and tables.
  • MathVista focuses on complex mathematical problem solving that integrates textual and visual information.
  • AI2D evaluates the ability to interpret scientific diagrams commonly used in elementary science education.
  • CapsBench emphasizes compositional reasoning by requiring models to generate comprehensive, detailed, and accurate descriptions of visual scenes. It challenges models to precisely capture a wide range of visual information, including object attributes, spatial relationships, and inter-object interactions.
Together, these benchmarks offer a comprehensive assessment of image-based multimodal capabilities.
We conduct evaluations on a diverse set of video benchmarks, including MMWorld [hemmworld], PerceptionTest [patraucean2023perception], Video-MME [videomme], MLVU [MLVU], MVBench [li2024mvbench], EventHallusion [zhang2024eventhallusion], TempCompass [tempcompass], VinoGround [vinoground], and DREAM-1K [wang2024tarsier].
  • MMWorld evaluates MLLMs’ ability to reason about real-world dynamics across diverse disciplines and tasks. It includes 1,910 videos and 6,627 QA pairs covering explanation, counterfactual reasoning, and future prediction.
  • PerceptionTest evaluates the perceptual and reasoning skills of MLLMs across video, audio, and text modalities. It includes 11.6K real-world videos and focuses on cognitive skills and reasoning types—such as memory, abstraction, and counterfactual thinking—beyond traditional classification or detection tasks. We use the validation set in the experiments.
  • Video-MME is a comprehensive benchmark for evaluating MLLMs across diverse video types, temporal lengths, and multimodal inputs including subtitles and audio. It features 900 manually annotated videos spanning 254 hours and 2,700 QA pairs, offering a rigorous test of models’ generalization and contextual understanding. We evaluate Video-MME without subtitles in our experiments.
  • MLVU is a benchmark designed for comprehensive evaluation of long video understanding, featuring extended video durations and diverse genres such as movies, surveillance, and egocentric videos. It includes a variety of tasks to assess MLLMs’ abilities in handling complex temporal dependencies and multi-scene reasoning across long-form content.
  • MVBench is a diagnostic benchmark designed to evaluate the temporal understanding capabilities of MLLMs through 20 challenging video tasks that go beyond static image reasoning. By systematically transforming static tasks into dynamic ones, it covers a wide range of temporal skills and ensures fair evaluation using ground-truth annotations converted into multiple-choice questions.
  • EventHallusion is a benchmark designed to evaluate hallucination in MLLMs, specifically focusing on event-level understanding—a core aspect of video analysis. It probes models’ susceptibility to language priors and vision-language biases, providing a targeted assessment of their reliability in temporal event reasoning.
  • TempCompass is a benchmark designed to evaluate the fine-grained temporal perception abilities of MLLMs across diverse task types. By introducing videos with controlled temporal variations and minimizing static or linguistic bias, it enables precise assessment of model performance on aspects such as speed, direction, and sequence understanding.
  • VinoGround is a benchmark that evaluates temporal counterfactual reasoning in short videos through 1,000 natural video-caption pairs.
  • DREAM-1K is a challenging benchmark for detailed video description, featuring 1,000 clips from diverse sources such as films, stock footage, and short-form videos. Each video is paired with fine-grained human-annotated descriptions, and evaluated using AutoDQ, a metric better suited for assessing rich, multi-event narratives than traditional captioning scores.
These benchmarks collectively cover a wide range of video understanding challenges, such as temporal reasoning, event prediction, visual grounding, perception under uncertainty, and multi-turn video-based instruction following, enabling a comprehensive assessment of the model’s performance across different video-centric tasks.
D Needle in a Haystack Test
Figure 11: Results of NIAH of Mavors with at most 60 video chunks.
[…]
E Showcases of Mavors in Image Captioning
We present a few examples of Mavors’ performance on the CapsBench benchmark in Figure 12. Mavors demonstrates a strong ability to generate accurate and detailed image captions, and it could recognize more proper nouns and understand human interactions and emotions more precisely than Qwen2.5-VL-7B.
Figure 12: Comparison of the generated image captions from Qwen2.5-VL-7B and Mavors-7B. The text in red contains wrong content, and the text in green marks the detailed descriptions only appear in Mavors.
F Showcases of Mavors with Token Compression
Our analysis indicates that as token compression increases up to 60%, there is negligible impact on Video QA performance, while performance on Captioning tasks degrades progressively. We present two case studies to illustrate the specific effects of captioning. In the first case (Figure 13), we observe that despite an imperfect initial caption, higher compression rates lead to increased model hallucinations. The model appears to perceive visual changes from the retained tokens but resorts to speculation, resulting in inaccurate descriptions. In the second case (Figure 14), increased compression causes the model to generate sparser descriptions, omitting critical details and introducing hallucinations. These findings suggest that token compression can pose performance risks, particularly for complex scene captioning tasks.
Figure 13: Example of captioning task with token compression: higher compression ratio leads to the missing of critical details.
Figure 14: Example of captioning task with token compression: higher compression ratio leads to the missing of critical details.